Find a quadratic equation for the solutions of 2 and 8
(x-2)(x-8) = 0

Answer by ewatrrr(24785)

 
Hi,
Find a quadratic equation for the solutions of 2 and 8
(x-2)(x-8) = 0
F First terms
O Outside terms
I Inside terms
L Last terms
x^2 - 10x + 16 = 0
y = x^2 -10x+16